# DynVFX: Augmenting Real Videos with Dynamic Content

### v1

@@ 4.3. Content Harmonization
$\mathcal{V}_{\text{orig}}$, $\mathcal{P}_{\text{VFX}}$triangleright Input video & instruction prompt
$\mathcal{\tau}_{\text{A}}$triangleright Extended Attention threshold
${\Psi}$triangleright Video segmentation model
@@
$\mathcal{P}_{\text{comp}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Composition Prompt
$\mathcal{O}_{\text{orig}},\mathcal{O}_{\text{edit}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Original objects and VFX object
${M}_{orig}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{V}_{\text{orig}},\mathcal{O}_{\text{orig}})$triangleright Extract source masks
${x}_{orig}\leftarrow\texttt{Encode}[\mathcal{V}_{\text{orig}}]$triangleright Encode video into latent space
@@
$\bm{x}_{res}=\bm{0}$triangleright initialize the residual latent
@@
else $\mathbb{\text{else }}{K^{E},V^{E}}\leftarrow\emptyset$
@@
$\mathcal{\hat{V}}_{comp}\leftarrow\texttt{Decode}(\bm{\hat{x}}_{comp})$triangleright Decode latent
$\bm{M}_{VFX}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{\hat{V}}_{comp},\mathcal{O}_{\text{edit}})$triangleright Extract VFX masks
@@
${\mathcal{V}}_{\text{comp}}\leftarrow\texttt{Decode}[{x}_{comp}]$triangleright Output video
@@
Figure 6. Qualitative comparison. Sample results of our method, SDEdit (Meng et al., 2022), DDIM inversion (Song et al., 2020), Lora fine-tuning (Hu et al., 2021), and Gen-3 (R Team, Runway, [n. d.]). See SM for videos.
5. Results
We evaluated our method on a dataset of 18 publicly sourced videos, featuring a wide range of complex scenes in terms of camera and object motion, lighting conditions, and physical environments. Additionally, we use some videos from DAVIS (Pont-Tuset et al., 2017). Our videos and implementation details are available in the SM.
Figures 15 show sample results of our method. As seen, our method facilitates natural integration of broad range of visual effects, ranging from environmental effects (tsunami in Fig. 4 and explosion in Fig. 5) to new object insertion (horse riding knight in Fig. 5 and dancing bear in Fig. 6). In all examples, the new content is naturally localized in the scene, even in challenging scenarios of multiple objects (dinosaurs or workers in Fig. 5) and partial occlusions (puppy in Fig. 1 and giraffe in Fig. 5).
5.1. Qualitative Evaluation
To the best of our knowledge, no existing method has been designed to synthesize dynamic objects in a given real video without user masks or any user information except for the input video itself and a simple instruction. We thus compare our method to the following baselines: (i) SDEdit (Meng et al., 2022) using the same T2V model as ours, (ii) DDIM inversion (Song et al., 2020) and sampling with the target prompt, (iii) LORA fine-tuning (Hu et al., 2021) of the T2V model and sampling with the target prompt and (iv) Gen-3 (R Team, Runway, [n. d.]) video-to-video, designed for video stylization.
Figure 6 shows a qualitative comparison to the baselines. As can be seen, all baselines exhibit different limitations in maintaining scene fidelity while introducing new content. SDEdit (Meng et al., 2022) manages to fulfill the edit prompt, yet the scene has significantly deviated from the original one in terms of appearance, motion, positioning, or scale (e.g., deer in the creek). DDIM inversion is not suitable for editing. LORA fine-tuning suffers from the trade-off between preserving aspects of the original scene and adding new content to the scene. Either over-fitting original scene appearance (e.g., added dragon-dog hybrid), or under-fitting the original layout (e.g., incorrect scale of deer). Gen-3 is conditioned on depth maps extracted from the input video, hence tends to significantly alter scene appearance and not allow the insertion of new objects that change the scene layout. In each case, these limitations affect the overall scene coherence and realism of the added elements. Our method successfully adds new content to the scene, achieving high fidelity to the user instructions while allowing for natural interactions between original and added elements (e.g. natural interaction between woman and bear).
Method Metrics VLM-based evaluation User Study
Gen-3 0.130 0.285 0.418 0.610 0.374 0.379 97.65 93.33
LORA finetuning 0.277 0.361 0.812 0.787 0.756 0.759 92.22 81.11
DDIM inv. sampling 0.184 0.444 0.535 0.699 0.528 0.529 99.20 98.67
SDEdit (0.9) 0.272 0.332 0.794 0.799 0.754 0.756 98.91 82.13
SDEdit (0.6) 0.111 0.567 0.510 0.704 0.513 0.504 97.69 96.76
w/o AnchorExtAttn 0.317 0.697 0.775 0.724 0.683 0.691 89.30 88.89
w/o Iterative Refinement 0.295 0.760 0.817 0.789 0.769 0.760 85.80 86.42
Ours 0.311 0.775 0.860 0.803 0.796 0.785 - -
Table 1. Quantitative Evaluation. We assess the quality of our method compared to several baselines.
5.2. Quantitative Evaluation
We numerically evaluate our results using the following metrics:
(i) Edit fidelity. Following previous works (e.g., (Hsu et al., 2024; Tewel et al., 2024)), we measure per-frame Directional CLIP Similarity (Radford et al., 2021; Gal et al., 2021) to assess the alignment between the change in the source and the target prompt, and the change between the source and edited frames.
(ii) Original content preservation. We evaluate how well the edited video preserves the original content outside the modified region. To this end, we segment the new content in the edited video using (Zhang et al., 2024) , and compute the masked Structural Similarity Index (SSIM) over the complementary regions to the edited ones.
(iii) VLM quality evaluation. We employ a Vision-Language Model (VLM) for expanding the per-frame metrics above in the following manner. We input to the VLM several frames from the edited videos and instruct it to evaluate four key aspects: how well the edit follows the text prompt (Text Alignment), the overall visual quality of the edited frames (Visual Quality), how well the new content is harmonized with the source frames (Edit Harmonization and the realism of the added object’s dynamics relative to the scene (Dynamics Score). For each aspect, the VLM outputs a score between 0 and 1, with higher scores indicating better performance. Our evaluation protocol is included in the SM.
Figure 7. Metrics. We measure CLIP Directional score (higher is better) and masked SSIM (higher is better). Our method demonstrates a better balance between these two metrics.
Figure  7 and Table 1 present the results of the described metrics on a set of 27 video-edit text pairs comprising 20 unique videos. As shown, our method outperforms the baselines in both SSIM and Directional CLIP metrics, demonstrating superior edit fidelity and maintaining higher structural similarity in the unedited regions. The VLM-based evaluation aligns with this assessment and additionally shows that our method produces videos that achieve better content integration and greater motion realism.
(iv) User study. We conducted a user study to evaluate the ability to integrate the new content while preserving the original video. Participants were shown the input video, a text description of the new content, our result, and a baseline output. They were asked two questions: “Which video better preserves the original footage while adding new content?” and “Which video better integrates the new content in a realistic and seamless way?”. In total, we collected 3240 user judgments from 120 users. As seen in Table 1, our method is consistently preferred over all baselines.
5.3. Ablations
We ablate key design choices of our method: anchor extended attention and iterative updates of the edit, by excluding each component from our framework.
As seen in Fig. 4(c), omitting AnchorExtAttn leads to new content being misaligned relative to the original scene - the added content is poorly integrated into the original scene. Applying only the first iteration of our method (w/o iterative refinement, Fig. 4(d)) results in a better alignment with the input video, but the composition is still unstable, as evident, for example, in the puppy’s body hovering over the box in the first scene. Our full method achieves better composition with proper spatial relationships, demonstrating the importance of both components for realistic scene editing Fig. 4(f). We numerically evaluate each ablation with the same set of metrics described in Sec 5.2 and report them in Table 1.
Figure 8. Limitations. In some cases, the T2V diffusion model struggles to precisely follow the edit prompt
6. Discussion and Conclusions
We introduced the task of augmenting real videos with new dynamic content based on a user-provided instruction. We presented a zero-shot method utilizing the T2V diffusion model in a feature manipulation framework, enabling correct localization and natural blending of new content with existing video elements.
As our method is built upon the pre-trained T2V diffusion model, the quality of the generated edits is inherently tied to the performance and capabilities of the underlying model. As seen in Fig. 8, the T2V model sometimes struggles with generating videos precisely following the edit prompt. Additionally, the text-based localization relies on the capabilities of the segmentation model (Zhang et al., 2024), which can sometimes produce inaccurate masks and fails to account for effects like shadows and reflections if not specified in the text prompt. Despite the limitations, our method significantly improves over baselines, expanding the capabilities of pre-trained text-to-video diffusion models.
7. Acknowledgement
This project received funding from the Israeli Science Foundation (grant 2303/20).
Appendix A Implementation Details
A.1. Models
Text-to-Video Model. We use a publicly available CogVideoX-5B (Hong et al., 2022; Yang et al., 2024) text-to-video model, which can generate videos with up to 480x720 pixel resolution, 6 seconds in length, 49 frames at 8 fps. This model is a transformer-based model that processes both text and video modalities together.
Segmentation Model. To segment the prominent objects in the video and the newly generated content, we utilize EVF-SAM2 (Zhang et al., 2024) - a text-based video segmentation model based on SAM2 (Ravi et al., 2024).
Visual Language Model.. Our vision-language model of choice is GPT-4o (Achiam et al., 2023), which we use through the official OpenAI API.
@@ A.2. Keys and Values Extraction
Following (Tumanyan et al., 2023; Yatim et al., 2023), to obtain T2V diffusion model intermediate latents, we use DDIM inversion (applying DDIM sampling in reverse order) on the input video, using 250 forward steps, with an empty string as text prompt. During the forward pass in our method, the intermediate latents are used for the extraction of keys and values.
A.3. Latent Mask Extraction
As discussed in Sec. 4.3, we iteratively update the residual latent $\bm{x}_{res}$ in the regions where the new content appears. This requires calculating the mask of the new content in the latent space. To do this, we first apply the segmentation model (Zhang et al., 2024) to the current output of SDEdit and get the mask of the new content in RGB space. However, the VAE in the T2V diffusion model involves both spatial and temporal downsampling, making it challenging to directly map RGB pixels to their corresponding latent regions. To address this, we encode the RGB masks through the VAE-Encoder and apply clustering to partition the resulting latents into two groups, effectively producing downsampled masks that align with the latent space representation.
A.4. Runtime
Our method’s two most computationally intensive parts are - DDIM inversion, which takes ~15 minutes, and iterative updates of the edit residual, which takes ~20 minutes. Importantly, DDIM inversion needs to be performed only once per video and can support multiple subsequent edits, making the process more efficient when applying various modifications to the same video content.
@@ Appendix B Baselines comparison details
For LORA fine-tuning baseline, we use the following default hyperparameters: Adam optimizer (Kingma and Ba, 2014), $41e-$ 4 learning rank, LORA rank 128, 800 fine-tuning steps. For comparison with GEN-3 (R Team, Runway, [n. d.]), we utilize the Gen-3 Alpha model via the publicly accessible web-based API, setting the ”Structure Transformation” hyperparameter to 5.
Appendix C Additional comparisons
We perform an additional qualitative comparison to MagicVFX (Guo et al., 2024). As can be seen in Fig. 10, MagicVFX struggles to remain faithful to the original scene and has lower visual quality compared to our method.
Figure 10. Comparison to MagicVFX. The result of MagicVFX the output differs significantly from the original video.
Figure 11. Output example for protocol
Figure 12. VLM instructions used for generating the textual descriptions.
Figure 13. VLM evaluation protocol
Appendix D VLM Prompting
@@
While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene. To overcome this, we ask the model to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly. In this conversation, GPT-4o will ”consult” with a VFX artist about how the new content should be integrated into the scene. Based on their discussion, it will be asked to provide a caption that describes how the added content fits into the scene. This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment. In this prompt, we also ask the VLM to provide a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$. The full prompt for the VLM is shown in Figure  12.
@@
In addition, as discussed in Sec. 5.2 we utilize the VLM for interpretable quality assessment. The full set of instructions for the VLM can be seen in Fig. 13.

### v2

@@ A.1. Models
$\mathcal{V}_{\text{orig}}$, $\mathcal{P}_{\text{VFX}}$ $\triangleright$ Input video & instruction prompt
$\mathcal{\tau}_{\text{A}}$ $\triangleright$ Extended Attention threshold
$\Psi$} $\triangleright$ Video segmentation model
@@
$\mathcal{P}_{\text{comp}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$ $\triangleright$ Composition Prompt
$\mathcal{O}_{\text{orig}},\mathcal{O}_{\text{edit}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$ $\triangleright$ Original objects and VFX object
${M}_{orig}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{V}_{\text{orig}},\mathcal{O}_{\text{orig}})$ $\triangleright$ Extract source masks
${x}_{orig}\leftarrow\texttt{Encode}[\mathcal{V}_{\text{orig}}]$ $\triangleright$ Encode video into latent space
@@
$\bm{x}_{res}=\bm{0}$ $\triangleright$ initialize the residual latent
@@
$\mathbb{\text{else }}{K^{E},V^{E}}\leftarrow\emptyset$
@@
$\mathcal{\hat{V}}_{comp}\leftarrow\texttt{Decode}(\bm{\hat{x}}_{comp})$ $\triangleright$ Decode latent
$\bm{M}_{VFX}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{\hat{V}}_{comp},\mathcal{O}_{\text{edit}})$ $\triangleright$ Extract VFX masks
@@
${\mathcal{V}}_{\text{comp}}\leftarrow\texttt{Decode}[{x}_{comp}]$ $\triangleright$ Output video
@@ A.2. Keys and Values Extraction
Following (Tumanyan et al., 2023; Yatim et al., 2024), to obtain the T2V diffusion model intermediate latents, we apply DDIM inversion (applying DDIM sampling in reverse order) on the input video. We perform inversion solely to extract the keys and values for AnchorExtAttn, while the sampling itself is initialized from a noisy $x_{comp}$. For the inversion process, we use 250 steps, with an empty string as the text prompt. We derive keys and values from the noisy latents produced by the inversion process (not from the reconstruction). During the sampling iterations in our method, we use these extracted keys and values for AnchorExtAttn. (Eq. 3) randomly samples a selection of anchor points, which are re-selected in each layer of the model. For all our results, we fix the hyperparameters and only adjust the dropout percentage in the AnchorExtAttn: 5% dropout for background and 70% for foreground in regular edits, while global edits use 10% for the foreground.
A.3. Obtaining $M_{orig}$
Figure 10. Qualitative Examples of $M_{orig}$. Each example includes the input video, user instruction, original masks, and corresponding result. In the top-right whale edit, our VLM protocol identifies the prominent elements to be coral and divers. In the top left pirate ships edit, the prominent elements are the cliffs. In the bottom left puppy edit, the woman and the ocean are identified. And in the bottom right bear edit, the woman and the couch are chosen.
To get $M_{\text{orig}}$, we ask a VLM to provide a list of foreground objects in the original video $\mathcal{O}_{\text{orig}}$ (typically the most spatially prominent elements within a scene) and then we obtain corresponding masks using EVF-SAM (Zhang et al., 2024a) - a text-based segmentation model.
Specifically, given the user instruction, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, our VLM protocol identifies the prominent elements from the original scene using the system prompt in Fig. 18. Next, each element (represented by words) is fed to EVF-SAM to produce masks, and their union forms $M_{orig}$, which is used in our AnchorExtAttn. Examples of such masks can be seen in Fig. 10.
A.4. Latent Mask Extraction
As discussed in Sec. 4.3, we iteratively update the residual latent $\bm{x}_{res}$ in the regions where the new content appears. This requires calculating the mask of the new content in the latent space. To do this, we first apply the segmentation model (Zhang et al., 2024a) to the current output of SDEdit and get the mask of the new content in RGB space. However, the VAE in the T2V diffusion model involves both spatial and temporal downsampling, making it challenging to directly map RGB pixels to their corresponding latent regions. To address this, we encode the RGB masks through the VAE-Encoder and apply clustering to partition the resulting latents into two groups, effectively producing downsampled masks that align with the latent space representation.
A.5. Runtime and Memory Usage
Our method’s most computationally intensive parts are - DDIM inversion, which takes ~15 minutes, and iterative updates of the edit residual, which takes ~20 minutes, while a single iteration of sampling with AnchoExtAttn takes between 2-5 minutes (depending on the noise level). Importantly, DDIM inversion needs to be performed only once per video and can support multiple subsequent edits, making the process more efficient when applying various modifications to the same video content. Querying the VLM and EVF-SAM adds negligible runtime (~5 and ~20 seconds correspondingly). We run our method on a NVIDIA A100, with 72GB memory in usage.
@@ Appendix B Baselines Comparison Details
The Baseline runtimes are (i) SDEdit - ~5 minutes (ii) DDIM inversion + Sampling - ~20 minutes (iii) LORA fine-tuning - ~2 hours, (iv) Gen-3 - ~30 seconds, (v) FlowEdit - ~3 minutes, (vi) AnyV2V - ~3 minutes and (vii) I2VEdit - ~30 minutes.
In terms of utilized Video Models, baselines (i) SDEdit, (ii) DDIM-inversion+Sampling, and (iii) LORA fine-tuning utilized the same text-to-video model (Yang et al., 2024) as our method, (iv) Gen-3 utilizes a proprietary video-to-video model by Runway (Alpha model via the publicly accessible web-based API), (v) FlowEdit is applied with Hunyuan Video model (Kong et al., 2024), (vi) AnyV2V uses I2VGen-XL model (Zhang et al., 2023b), supporting 15 frames, and (vii) I2VEdit utilizes Stable Video Diffusion model (Blattmann et al., 2023) supporting 27 frames. For AnyV2V and I2VEdit, the first frames were edited with Gemini image editing (Gemini 2.0, 2024).
@@
For LORA fine-tuning baseline, we use the following default hyperparameters: Adam optimizer (Kingma and Ba, 2014), 1e41e-4 learning rank, LORA rank 128, 800 fine-tuning steps. For comparison with Gen-3 (Runway, 2024), we set the ”Structure Transformation” hyperparameter to 5.
For FlowEdit, we use source and target prompts produced from our VLM-VFX protocol, as this method heavily relies on source and target prompting.
Appendix C Limitations
Figure 11. Failure Case Example. When trying to add content that already exists in the scene (e.g., an additional dog), it is best to specify the differences between them in the user instructions. As can be seen, when trying to add a Poodle (similar appearance and size to a Havanese), EVF-SAM fails to distinguish between the two. However, when trying to add a German Shepherd, the segmentation model manages to make the distinction.
Figure 12. Comparison to Reference-Based Insertion Methods. The left plane shows the results of our method, while the right plane shows the results of Pika and Kling. As can be seen, these reference-based methods struggle to generate global effects.
Figure 13. Comparison to MagicVFX. MagicVFX (second column) deviates significantly from the original video. As seen in the third column, our method successfully adds new content while maintaining high fidelity to the original scene.
Our method relies on the target segmentation provided by a text-based segmentation model (Zhang et al., 2024a). While this solution often succeeds for most edits, it can sometimes produce inaccurate masks and fail to account for effects like shadows and reflections if not specified in the text prompt. To handle such cases, users can add a specification in text to include these secondary effects. Furthermore, the segmentation model may struggle in scenarios where multiple entities with similar visual characteristics are present in the scene. As shown in Fig.  11, segmentation issues can occur when adding content of similar appearance and size. However, when the entities are distinguishable enough, the new content is successfully integrated into the scene.
Furthermore, our framework performs a single-shot edit and does not support interactive, in-context refinements - once an edit is applied, users cannot request follow-up adjustments (e.g., tweaking object placement or appearance) without restarting the process. While VFX pipelines typically involve several feedback-driven iterations to perfect a scene, our current design does not support such incremental revisions. Future work could integrate an interactive editing loop, potentially via an autoregressive T2V model, to allow users to make successive edits without reinitializing the entire workflow.
Figure 14. Additional Example for Ablations. (b) Excluding both AnchorExtAttn and the Iterative refinement process results in significant misalignment with the original scene and poor harmonization (e.g., the size of the puppy relative to the scene and boundary artifacts). (c) Omitting AnchorExtAttn leads to incorrect positioning of the new content. (d) Removing iterative refinement results in poor harmonization. Our full method (e) exhibits good localization and harmonization of the edit.
Appendix D Additional Qualitative Comparisons
We perform additional qualitative comparisons to MagicVFX (Guo et al., 2024) and two reference-based methods: Pikadditions (Pika, 2024) and Kling (KlingAI, 2024). As can be seen in Fig. 13, MagicVFX struggles to remain faithful to the original scene and has lower visual quality compared to our method. Furthermore, as illustrated in Fig. 12, both Pikadditions and Kling struggle to generate global effects. As can be seen, the audience members are incorrectly overlaid in front of the dancer. Additionally, the dinosaur occludes one of the bikers, and only one dinosaur is generated, although we prompt for a global edit. This demonstrates scenarios where text-based insertion is preferable over reference-based insertion. Even though such methods do not explicitly require the control of location and dynamics of the new content (as our method), they rely on the reference assets, which limits their applicability whenever a suitable example cannot be found for the given video.
Appendix E VLM Prompting
E.1. Our VLM as a VFX Assistant Protocol
@@
Although the model is capable of providing descriptive captions of the source scene, in some cases, we observed that it fails to provide captions suitable for compositing VFX with the scene. To overcome this, we ask the model to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly. In this conversation, GPT-4o will ”consult” with a VFX artist about how the new content should be integrated into the scene. Based on their discussion, it will be asked to provide a caption that describes how the added content fits into the scene. This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment. In this prompt, we also ask the VLM to provide a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$. The full prompt for the VLM is shown in Fig.  17.
@@
In addition, as discussed in Sec. 5.2, we utilize the VLM for interpretable quality assessment. The full set of instructions for the VLM can be seen in Fig. 19.
E.2. Ablation of the VLM as a VFX Assistant Protocol
Figure 15. Ablating VLM Protocol. Without the protocol (middle), the VLM often yields an unsuitable prompt for adding the desired content, while our protocol (right), gives $\mathcal{P}_{\text{comp}}$ which allows for scene-aware additions to $\mathcal{O}_{\text{orig}}$ (see jellyfish and parrot behind the window). Additionally, as can be seen in Table 1 in the main paper, Table 2 and Fig. 9 from the SM, the metrics for this ablation are worse than those for our full method.
To validate the importance of our VLM protocol, we performed an ablation by prompting the VLM with a simplified system prompt, asking it to caption the original video and provide the edit prompt. As seen in Fig. 15, simplifying the system prompt results in misinterpretation of the instruction with respect to the scene or fails to add new content in general.
Appendix F VLM as Evaluation Metric Validation
Table 2. VLM Quality Evaluation with an alternative backbone (Claude Haiku 3.5). We report text alignment, visual quality, edit harmonization, and dynamics score. The results closely match those obtained with GPT-4o in the main paper, showing consistent outcomes across all metrics and methods.
Method Text Align. Visual Quality Edit Harm. Dynamics Score
Gen-3 0.517 0.499 0.412 0.445
LORA finetuning 0.807 0.755 0.722 0.719
DDIM inv. sampling 0.423 0.487 0.384 0.385
SDEdit (0.9) 0.769 0.734 0.697 0.682
SDEdit (0.6) 0.420 0.561 0.425 0.405
FlowEdit 0.670 0.651 0.591 0.605
AnyV2V 0.693 0.683 0.625 0.625
I2VEdit 0.769 0.747 0.697 0.698
w/o AnchorExtAttn 0.656 0.620 0.547 0.556
w/o Iterative Refinement 0.704 0.672 0.631 0.634
w/o VLM Protocol 0.697 0.700 0.631 0.644
Ours 0.820 0.760 0.721 0.721
To show that our VLM-based evaluation aligns with human preferences, we performed the following assessments: Measured Pearson’s correlation coefficients between our VLM scores (Edit Harmonization and Dynamics) and the user study scores for Edit Harmonization, as both assess content integration quality. We observe a strong correlation $r=0.82$ ($p=0.025$) and $r=0.89$ ($p=0.007$). We compute Pearson’s correlation between the Text Alignment score and the CLIP Directional score - a metric commonly used to assess text alignment and shown to correlate with human preference (Kim et al., 2025) - and found a strong correlation: r=0.81 (p=0.029). These strong correlations suggest that our metric aligns with human judgments in the context of our task.
We utilize the same VLM (GPT-4o (Achiam et al., 2023)) in our pipeline and evaluation. However, this does not introduce bias in our evaluation due to the separation between the internal representations of the T2V model and the VLM’s reasoning. During evaluation, the VLM does not “recognize” the video generated by the T2V model as its own, making the assessment unbiased and independent of its earlier role. We further validate this by evaluating the results with a different VLM (Claude Haiku 3.5 (Anthropic, 2024)) in Table 2 and demonstrate that it provides similar scoring to that reported in the paper.
Our VLM-based evaluation metrics are inspired by recent work such as AutoVFX (Hsu et al., 2025) and (Wu et al., 2024), which demonstrate the effectiveness of VLMs for automatic quality assessment tasks.
Figure 16. VLM as a VFX Assistant Overview. Given instructions and an input video, the VLM (grey) interprets the interaction, reasons about the integration, and composes $\mathcal{P}_{\text{comp}}$ to guide the T2V generation. Inputs include: (1) a system prompt (red) guiding the VLM to produce VFX reasoning (purple), (2) a user prompt describing the content to integrate into the scene (green), and (3) key frames from the input video. The VLM produces VFX reasoning (purple) describing spatial placement, lighting, and motion for seamless integration, followed by $\mathcal{P}_{\text{comp}}$ (blue) captioning the target scene. Next, given the user prompt, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, the VLM identifies prominent elements from the original scene $\mathcal{O}_{\text{orig}}$ (yellow) and the new content to be added $\mathcal{O}_{\text{edit}}$ (brown). These outputs are used in our framework for the application of SAM, used to aid localization and harmonization of the new content. As can be seen in the output video, in this example, a majestic elephant is seamlessly integrated into the scene.
Figure 17. System Prompt We Input the Pre-trained VLM (VFX Assistant). The system prompt contains guidelines to supply the VLM with the context of our task in order for it to interpret the user instruction, reason about the interaction, and accordingly, output a scene description that will be used for editing.
Figure 18. System Prompt For Obtaining $\mathcal{O}_{\text{orig}}$ and $\mathcal{O}_{\text{edit}}$. Given the user prompt, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, the VLM identifies prominent elements from the original scene $\mathcal{O}_{\text{orig}}$ and the new content to be added $\mathcal{O}_{\text{edit}}$. These outputs are used in our framework for the application of SAM, used to aid localization and harmonization of the new content.
Figure 19. System Prompt for VLM-Based Evaluation. containing the guidelines for the evaluation setup, where original and edited frames are compared against an edit prompt across four criteria: prompt alignment, visual quality, content harmonization, and dynamics.